You've set your hypotheses on the mean monthly rent of all advertised one-bedroom apartment as follows.

Test it at α = 0.05 (Sample size 100). Please write your conclusion.

  • Ho: µ = 550, Ha: µ ≠ 550
Answer #1

Solution :

This is the two tailed test .

The null and alternative hypothesis is ,

H0 :  \mu  = 550

Ha : \mu    \neq 550

Test statistic = t

n =100

df = n- 1 = 100-1 = 99

Test statistic = t  = 1.984

P-value =0.0500

\alpha = 0.05  

P-value =\alpha

0.0500 = 0.05

Accepets the null hypothesis .